OpenStudy (anonymous):

use the formula A=P(1+rt) to find the indicated quantity. allan borrowed $4700 from his father to buy a car. he repaid him after 9 months with interest of 4% per year. find the total amount he repaid....... A. 141 B. 4888 C. 4825.33 D. 4841

OpenStudy (anonymous):

i think its c but not too sure

OpenStudy (ivettef365):

A = 4700(1 + (0.04 × 0.75)) = 4841

OpenStudy (uri):

Where did 0.75 come from?

OpenStudy (ivettef365):

because 9 month represents .75 of a year

OpenStudy (ivettef365):

9/12 = .75

OpenStudy (uri):

Oh alright and P stands for?

OpenStudy (ivettef365):

P = Principal

OpenStudy (anonymous):

thank you everyone!!!!!

OpenStudy (uri):

I knew it...And A=Amount...so oh got ya what mistake i did

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

can you help me with one more question please its to figure out how many months for sometihng

OpenStudy (uri):

One more thing @ivettef365 Where did the r go? r=Rate..you just wrote 0.75 that is t.

OpenStudy (ivettef365):

rate = 4% = 0.04 that is what I multiply by 0.75

OpenStudy (ivettef365):

(1+rt) which is 1 + (rate x time)

OpenStudy (uri):

Got it,sorry to annoy :P

OpenStudy (ivettef365):

not a problem at all :)

OpenStudy (anonymous):

Try out simple interest calculator https://www.easycalculation.com/simple-interest.php